OpenStudy (anonymous):

36 = 5/9k. solve for k.

OpenStudy (anonymous):

I will award medals.

OpenStudy (anonymous):

k=5/324

OpenStudy (anonymous):

The possable answers are. A. 4/5 B. 20 C. 64/45 D. 180

OpenStudy (mertsj):

\[36=\frac{5}{9}k\] \[\frac{9}{5} \times 36=\frac{9}{5}\times \frac{5}{9}k\]

OpenStudy (anonymous):

so c?

OpenStudy (mertsj):

None of those answers work.

OpenStudy (mertsj):

is it 5/(9k) or (5/9)k

OpenStudy (anonymous):

I'm on a practice test and these are the answers they gave me.

OpenStudy (mertsj):

Please answer my question.

OpenStudy (anonymous):

I think It's ( 5/9 )k

OpenStudy (anonymous):

..thats weird did you type it right?

OpenStudy (mertsj):

Well then, the answer is 64.8 or 64 4/5

OpenStudy (anonymous):

I copied and pasted it.

OpenStudy (mertsj):

\[64\frac{4}{5}\]

OpenStudy (anonymous):

yeah because 324/5 = c if c is 64 4/5 not 64/45